if erin 3 times warrens age. and in 4 years she'll be twice the age of warren. how old are each of them now?: if erin 3 times warrens age. and in 4 years she'll be twice the age of warren. how old are each of them now?
Answer by mangopeeler07(423)

erin is 3 times warrens age:

warren=w
erin=3w

in 4 years she'll be twice the age of warren:

erin+4=2(warren+4)
3w+4=2(w+4)

how old are each of them now?
dsitribute--
3w+4=2w+8

Subtract 2w--
w+4=8

Subtract 4--
w=4

Therefore,
**Warren=4
**Erin=12

Check: 3w+4=2(w+4) becomes 12+4=2(8) or 16=16. So the ages are right.